# Zero-Shot Hyperspectral Pansharpening Using Hysteresis-Based Tuning for Spectral Quality Control

### v1

@@ I Introduction
Hyperspectral (HS) remote sensing imagery is a key resource for monitoring and managing such diverse environments as forests [1, 2], agricultural areas [3, 4], oceans [5, 6], coastal areas [7, 8], inland waters and wetlands [9, 10], snow and ice [11, 12], atmosphere [13, 14], mineral resources [15, 16] [17], urban areas [18, 19]. Due to the technological constraints of the sensors and the very high number of spectral bands acquired, HS images have a rather low spatial resolution. A cost-effective solution to overcome this limitation is to equip the flight system with an additional camera to acquire, in synchrony with the HS image, a single high-resolution panchromatic (PAN) band. Then, using a fusion process known as pansharpening the PAN-HS pair is merged into a single high-resolution HS datacube. This data fusion technique enhances spatial resolution while preserving spectral consistency. Algorithms designed to improve the spatial resolution of remotely sensed images have already been shown to significantly boost performance in downstream tasks such as classification [20] and land use/land cover mapping [21]. It is therefore reasonable to expect that enhanced HS data could offer substantial benefits in applications where HS information is crucial but often underexploited due to limited spatial resolution, for example, in air pollution estimation [22] or water quality monitoring [23]. Examples of satellite-borne missions that feature the joint acquisition of PAN and HS images are PRISMA (PRecursore IperSpettrale della Missione Applicativa) by the Italian Space Agency (ASI), and EO-1/ALI (Earth Observing-1/Advanced Land Imaging) by the U.S. National Aeronautics and Space Administration (NASA). In both cases, the spatial resolution is 30m for the HS component and much better for the PAN image, 5m for PRISMA and 10m for EO-1/ALI. Needless to say, the PRISMA case is certainly more interesting, as it promises to provide high-resolution data, but also more challenging since it requires the pansharpening algorithm to handle a resolution ratio of 6 between the available HS data in input and the desired output.
@@ III-A Correlation-based Spatial Loss for HS Images
In our proposal, we keep the two-component loss formulation (1) previously adopted [73] in unsupervised pansharpening and use the same spectral component of [75], based on the $\ell_{1}$-norm (2), which shows good convergence properties and outperforms other norms, such as $\ell_{2}$ or ERGAS, in terms of perceptual quality. In any case, the exact form of the spectral loss is not really critical as it is always obtained by comparing the pansharpened product $\widehat{\mathbf{\mathbf{H}}}$ with a clear reference, the interpolated HS image $\widetilde{\mathbf{\mathbf{H}}}$ or (after decimation) the low-resolution HS itself. We focus instead on the spatial loss term, which is more elusive as it is intimately related to the problem of assessing spatial quality without a reference. In R-PNN we used the correlation-based loss (3) originally proposed in [72] for MS pansharpening. The idea is to push the pansharpened band to have a high local correlation with the PAN, so as to exhibit the same spatial layout. The correlation coefficient measures how well one image can be linearly predicted from another. A coefficient of $|\rho|=1$ indicates perfect predictability, while $\rho=0$ indicates no correlation. Ideally, pansharpened bands should closely match the spatial structure of the PAN image, resulting in high correlation. However, to preserve spectral information, this correlation cannot be unitary. In fact, it is expected to vary spatially and among different bands, depending on the scene’s characteristics. For this reason, to prevent the injection of alien details from the PAN into the spectral band, the correlation itself is capped by its estimate, $\rho^{\rm max}(s)$, computed on reduced resolution data.
This correlation-based loss proved extremely effective for MS pansharpening, ensuring results characterized by very high spatial fidelity [72, 73]. Its extension to the HS case, however, is more controversial. In fact, the basic assumption is that $\rho(s)$ is generally high and certainly positive. While this is more than reasonable for the MS case, it becomes largely questionable in the HS case because many spectral bands fall outside the visible range (i.e., the PAN range) and exhibit a lower correlation with it, and sometimes the so-called inversion phenomenon [77] with regions having strong but negative correlation. An algorithm that disregards these physical principles may produce pansharpened results with visible artifacts, such as unnatural color variations and exaggerated contrasts. Furthermore, during the target adaptation phase, this inconsistency could lead to a higher conflictual behavior between the spectral and spatial loss functions, ultimately hindering the overall reconstruction quality.
@@
A few observations are in order. The sample band from the visible spectrum shows a very high correlation with the PAN, 0.95 on the average and never negative. This is a common behavior in the visible spectrum, more pronounced in the green range and a little less towards the coastal and red/infrared limits. In this situation, maximizing the local correlation, without any upper bound, makes perfect sense. Outside the visible spectrum, however, the correlation field has a much higher dynamics, and our sample band has an average correlation of just 0.33 with the PAN. Such a low value, however, does not signal lack of dependence. The spatial layouts keep being mostly the same, as easily appreciated in the close-up image, but several regions exhibit a strong negative correlation. This is the case, for example, of the circularly shaped elements clearly visible in the close-up, which are very dark in the PAN, and very bright in $\mathbf{H}_{o}$. In addition, intermediate correlation values, between -0.6 and 0.6, occur at the boundaries between positively and negatively correlated regions just because of the finite size of the estimation window (6$\times$6 in our experiment). This latter problem impacts also on the estimation of the bounding map, $\rho^{\rm max}$, further reducing its effective resolution and making it scarcely reliable.
@@
Note that we have also removed the bounding map $\rho^{\rm max}$ due to its limited reliability. However, we will show that a suitable hysteresis-based optimization schedule will prevent the injection of inappropriate PAN details in the pansharpened image.
@@ IV-A2 Limiting the Number of Iterations
$N^{\rm max}_{b}=N_{0}+\Delta N(1-c_{b})$ (8)
with $c_{b}$ the correlation coefficient between bands $b$ and $b-1$. Therefore, the maximum number of iterations is $N_{0}\geq N_{S}^{\rm max}$ in case of perfect correlation ($c_{b}=1$) and $N_{0}+\Delta N$ in case of orthogonal bands ($c_{b}=0$). With this rule, we can easily set an overall bound to the total number of iterations $N$ as follows
@@
$\Delta N=\frac{\eta BN_{0}}{B-\sum_{b}c_{b}}$ (10)
Fig. 9 shows an example of how iterations distribute over the various bands of a PRISMA image. Here we set $N_{S}^{\rm max}=20,N_{0}=50$ and $\eta=8$. $N_{b|S}$ is shown in yellow and $N_{b}$ in brown, with its upper bound $N^{\rm max}_{b}$ in light gray. Although a maximum of $(1+\eta)N_{0}=450$ iterations per band is allowed on average, only 47 are used on the average. Most of the times, tuning stops because $N_{b|S}$ reaches its maximum, $N_{S}^{\rm max}=20$. The total number of iterations rarely exceeds $N_{0}=50$ and only for a group of bands, on the boundary between the visible and NIR ranges, the $N_{b}=N_{b}^{\rm max}$ exit condition is activated. The entire pansharpening procedure of the proposed method is summarized in Algorithm 1.
@@
Algorithm 1 Pseudo-code $\rho$-PNN
Hyperparameters =$=\{\gamma_{high},\gamma_{low},\beta_{0},\eta,N_{0},N_{S}^{max},\epsilon,\alpha\}$
$B\leftarrow size(H,3)$triangleright # of bands
$c\leftarrow correlations(H)$triangleright vector $c_{b}=corrcoef(H_{b},H_{b-1})$; $c_{0}=0$
$\phi\leftarrow weightInit(net)$triangleright random initialized net parameters
for $b\leftarrow 1,B$ do
    $N_{b}^{max}\leftarrow iterationUpperBound(c,N_{0},B,\eta)$triangleright eqs 8 and 10
    $\tilde{H}_{b}\leftarrow expInterpolator(H_{b})$
    $\mathcal{L}_{\lambda}^{EXP}\leftarrow l_{1}\text{-}norm(H_{b},subsample(\tilde{H}_{b}))$triangleright reprojection error on $\tilde{H}_{b}$
    $N_{b},N_{b|s}\leftarrow 0$triangleright counters initializations
    $\beta\leftarrow warmUp(H_{b},P,net,\phi,\beta_{0},\epsilon,\alpha)$triangleright warmup to fix $\beta$
    $full\leftarrow{{}^{\prime}OFF^{\prime}}$triangleright full loss (’ON’) vs spectral-only loss (’OFF’)
    while  $\{N_{b|S}<N_{S}^{max}\}\;\;\mbox{OR}\;\;\{N_{b}<N^{\rm max}_{b}$}  do
        $\phi,\mathcal{L}_{\lambda},\mathcal{L}_{S}\leftarrow tuningStep(H_{b},P,net,\phi,\alpha,\beta,full)$
        $N_{b}\leftarrow N_{b}+1$
        if {full} then
           $N_{b|S}\leftarrow N_{b|S}+1$
           if $\mathcal{L}_{\lambda}/\mathcal{L}_{\lambda}^{EXP}>\gamma_{high}$then
               $full\leftarrow^{\prime}OFF^{\prime}$triangleright swith-off
           else if $\mathcal{L}_{\lambda}/\mathcal{L}_{\lambda}^{EXP}<\gamma_{low}$then
               $full\leftarrow{{}^{\prime}ON^{\prime}}$triangleright swith-on
           end if
        end if
    end while
    $\hat{H}_{b}\leftarrow predict(H_{b},P,net,\phi)$triangleright inference on band $b$
end for
return $\hat{H}$
@@
$\gamma_{{}_{\rm HIGH}}$ 0.65 Superior limit fixed for ${\cal L}_{\lambda/E}$
$\gamma_{{}_{\rm LOW}}$ 0.59 Inferior limit for ${\cal L}_{\lambda/E}$
$\beta_{0}$ 2 Starting $\beta$, weight of the spatial loss term
$\epsilon$ 0.007 Tolerated growth factor of ${\cal L}_{\lambda}$ for $\beta$ adaptation
$N_{S}^{\rm max}$ 20 Target # of tuning iterations with active spatial loss
$N_{0}$ 80 Max # of iterations for bands with perfect correlation with the preceding one
$\eta$ 30 Overhead factor to globally constrain extra iterations
@@ IV-B Assessment Methodology
To assess the quality of the proposed solution we take advantage of the Benchmarking toolbox [42] which provides a publicly available set of rich and diverse PAN-HS PRISMA images, a large set of SotA comparative methods, and several performance evaluation tools based on the most credited assessment indexes. PRISMA was chosen for this evaluation since it provides a real panchromatic image acquired through a dedicated sensor operating in the 400 - 700 nm spectral range. In addition to the PAN sensor, PRISMA is equipped with two hyperspectral instruments: a VNIR sensor covering wavelengths from 400 to 1010 nm with 66 bands, and a SWIR sensor spanning 920 to 2505 nm with 173 bands, resulting in a total of 239 spectral channels. The dataset released by [42] consists of 16 PRISMA images, 12 designated for training and validation, and 4 for testing. To enhance data reliability, noisy spectral bands are removed, reducing the dataset to 159 clean and informative bands for benchmarking purposes. Additionally, to further investigate the generalizability of the proposed method, we considered an additional test dataset in which the PAN image was simulated using the RGB bands of the Copernicus Sentinel-2 mission (665,560 and 490 nm with 10m spatial resolution) and subsequently co-registered with the PRISMA HS data. Tab. IV gathers the comparative methods which can be roughly grouped in four categories: component substitution (CS), multiresolution analysis (MRA), model-based optimization (MBO), and deep learning (DL). The interested Reader is referred to [42] for details on these methods.
TABLE IV: HS pansharpening methods.
@@
The assessment follows the well-established Wald’s protocol [84] which considers two contexts: a reduced resolution framework, where suitably degraded (in resolution) datasets are used for assessment with GT (where the ground truth is given by the original full resolution HS image), and a full resolution framework for assessment without GT but using the original data as input. Accordingly, the quality indexes, gathered in Tab. V, are split in two groups for reduced resolution (RR) and full resolution (FR) assessment.
@@
TABLE V: HS pansharpening quality assessment indexes.
@@
TABLE VI: PAN-MS PRISMA benchmark datasets for testing.
Date Location Land covers
05/09/22 Cagliari, Italy Roofs, Streets, Crops, Hills, Water
24/08/23 Udine, Italy Roofs, Streets, Crops, Water
08/09/23 Ford County, Kansas Streets, Crops
20/11/23 Macuspana, Tabasco Streets, Crops
Real FR test image size: 1200$\times$1200
Resolution ratio $R=6$
Bands $B=159/239$
Resolution: 30 m (RR) or 5 m (FR)
Experiments are carried out on the test datasets of the Toolbox, summarized in Tab. VI, comprising four acquisitions/scenes. From each original PRISMA PAN-HS pair (PAN/HS size: 3600$\times$3600/600$\times$600) two test datasets were created: a FR one (PAN/HS size: 1200$\times$1200/200$\times$200) obtained by cropping; a RR one (PAN/HS/GT size 600$\times$600/100$\times$100/600$\times$600) obtained by resolution downgrading of the whole PAN-HS pair, with the original FR HS playing the role of GT. Regarding the additional test dataset with the Sentinel-2 simulated PAN, we selected a Sentinel-2 acquisition covering the same area as the PRISMA Kansas dataset (in Tab. VI), acquired with a one-day temporal shift, and co-registered it accordingly. To preserve the original PRISMA ratio of 6, the hyperspectral data were then downscaled by a factor of 2, resulting in a spatial resolution of 60 meters (PAN/HS size: 600$\times$600/100$\times$100).

### v2

@@ I Introduction
Hyperspectral (HS) remote sensing imagery is a key resource for monitoring and managing such diverse environments as forests [1, 2], agricultural areas [3, 4], oceans [5, 6], coastal areas [7, 8], inland waters and wetlands [9, 10], snow and ice [11, 12], atmosphere [13, 14], mineral resources [15, 16, 17], urban areas [18, 19]. Due to the technological constraints of the sensors and the very high number of spectral bands acquired, HS images have a rather low spatial resolution. A cost-effective solution to overcome this limitation is to equip the flight system with an additional camera to acquire, in synchrony with the HS image, a single high-resolution panchromatic (PAN) band. Then, using a fusion process known as pansharpening the PAN-HS pair is merged into a single high-resolution HS datacube. This data fusion process improves spatial resolution while preserving spectral coherence, thereby boosting the performance of downstream tasks such as classification [20] and land use/cover mapping [21]. Examples of satellite-borne missions that feature the joint acquisition of PAN and HS images are PRISMA (PRecursore IperSpettrale della Missione Applicativa) by the Italian Space Agency (ASI), and EO-1/ALI (Earth Observing-1/Advanced Land Imaging) by the U.S. National Aeronautics and Space Administration (NASA). In both cases, the spatial resolution is 30m for the HS component and much better for the PAN image, 5m for PRISMA and 10m for EO-1/ALI. Needless to say, the PRISMA case is certainly more interesting, as it promises to provide high-resolution data, but also more challenging since it requires the pansharpening algorithm to handle a resolution ratio of 6 between the available HS data in input and the desired output.
@@ III-A Correlation-based Spatial Loss for HS Images
In our proposal, we keep the two-component loss formulation (1) previously adopted [71] in unsupervised pansharpening and use the same spectral component of [73], based on the $\ell_{1}$-norm (2), which shows good convergence properties and outperforms other norms, such as $\ell_{2}$ or ERGAS, in terms of perceptual quality. In any case, the exact form of the spectral loss is not really critical as it is always obtained by comparing the pansharpened product $\widehat{\mathbf{\mathbf{H}}}$ with a clear reference, the interpolated HS image $\widetilde{\mathbf{\mathbf{H}}}$ or (after decimation) the low-resolution HS itself. We focus instead on the spatial loss term, which is more elusive as it is intimately related to the problem of assessing spatial quality without a reference. In R-PNN we used the correlation-based loss (3) originally proposed in [70] for MS pansharpening. The idea is to push the pansharpened band to have a high local correlation with the PAN, so as to exhibit the same spatial layout. Indeed, the correlation coefficient measures how well one image can be linearly predicted from another. Since pansharpened bands and PAN share the same spatial structure, they can be expected to be strongly correlated. On the other hand, each band has unique spectral properties that must be carefully preserved. Therefore, to prevent the injection of alien details from the PAN into the spectral band, the correlation itself, $\rho$, is capped by its estimate, $\rho^{\rm max}(s)$, computed on reduced resolution data.
This correlation-based loss proved extremely effective for MS pansharpening, ensuring results characterized by very high spatial fidelity [70, 71]. Its extension to the HS case, however, is more controversial. In fact, the basic assumption is that $\rho(s)$ is generally high and certainly positive. While this is more than reasonable for the MS case, it becomes largely questionable in the HS case because many spectral bands fall outside the visible range (i.e., the PAN range) and exhibit a lower correlation with it, and sometimes the so-called inversion phenomenon [75] with regions having strong but negative correlation.
@@
A few observations are in order. The sample band from the visible spectrum shows a very high correlation with the PAN, 0.95 on the average and never negative. This is a common behavior in the visible spectrum, more pronounced in the green range and a little less towards the coastal and red/infrared limits. In this situation, maximizing the local correlation, without any upper bound, makes perfect sense. Outside the visible spectrum, however, the correlation field has a much higher dynamics, and our sample band has an average correlation of just 0.33 with the PAN. Such a low value, however, does not signal lack of dependence. The spatial layouts keep being mostly the same, as easily appreciated in the close-up image, but several regions exhibit a strong negative correlation. This is the case, for example, of the circularly shaped elements clearly visible in the close-up, which are very dark in the PAN, and very bright in $\mathbf{H}_{o}$. In addition, intermediate correlation values, between -0.6 and 0.6, occur at the boundaries between positively and negatively correlated regions just because of the finite size of the estimation window (6$\times$6 in our experiment). This latter problem impacts also on the estimation of the bounding map, $\rho^{\rm max}$, further reducing its effective resolution and making it scarcely reliable. An algorithm that disregards these phenomena may generate visible pansharpening artifacts, such as unnatural color variations and exaggerated contrasts. Moreover, conflicts may arise between the spectral and spatial terms of the loss, hindering the training or fine-tuning phases.
@@
Note that we have also removed the bounding map $\rho^{\rm max}$ due to its limited reliability. In fact, a suitable hysteresis-based optimization schedule will prevent the injection of inappropriate PAN details in the pansharpened image.
@@ III-C Hysteresis-based Training Procedure
Algorithm 1 describes the proposed method more formally by means of a high-level pseudo-code.
Input: $\mathbf{P}$, $\mathbf{H}$, net (network structure)
Hyperparameters: $\gamma_{{}_{\rm HIGH}}$, $\gamma_{{}_{\rm LOW}}$, $\beta_{0}$, $\eta$, $N_{0}$, $N_{S}^{\rm max}$, $\epsilon$
Output: $\widehat{\mathbf{\mathbf{H}}}$
1:$B\leftarrow{\rm size}(\mathbf{H},3)$triangleright # of bands
2:$\mathbf{c}\leftarrow{\rm correlations}(\mathbf{H})$triangleright $\mathbf{c}_{b}={\rm corrcoef}(\mathbf{H}_{b},\mathbf{H}_{b-1})$; $\mathbf{c}_{1}=0$
3:$\phi\leftarrow{\rm weightInit}({\rm net})$triangleright net weights init.
4:for $b\leftarrow 1:B$ do
5:    $N_{b}^{\rm max}\leftarrow{\rm iterationUpperBound}(\mathbf{c},N_{0},B,\eta)$triangleright Eqs. (8), (10)
6:    $\widetilde{\mathbf{H}}_{b}\leftarrow{\rm interpolator}(\mathbf{H}_{b})$
7:    $\mathcal{L}_{\lambda}^{\rm EXP}\leftarrow{\rm getSpectralLoss}(\widetilde{\mathbf{\mathbf{H}}}_{b},\mathbf{H}_{b})$
8:    $N_{b},N_{b|S}\leftarrow 0$triangleright counter init.
9:    $\beta\leftarrow{\rm warmUp}(\mathbf{H}_{b},\mathbf{P},{\rm net},\phi,\beta_{0},\epsilon,\alpha)$triangleright $\beta$ warm-up
10:    ${\rm ON}\leftarrow\mathbf{\rm false}$triangleright ${\cal L}_{S}$ flag init.
11:    while  $\{N_{b|S}<N_{S}^{\rm max}\}\;\mbox{\bf and}\;\{N_{b}<N^{\rm max}_{b}$}  do $\triangleright$ Eq. (7)
12:        $\phi,\mathcal{L}_{\lambda},\mathcal{L}_{S}\leftarrow{\rm optimizationStep}(\mathbf{H}_{b},\mathbf{P},{\rm net},\phi,\beta,{\rm ON})$
13:        $N_{b}\leftarrow N_{b}+1$
14:        if ON then
15:           $N_{b|S}\leftarrow N_{b|S}+1$
16:        end if
17:        if ${\cal L}_{\lambda/E}>\gamma_{{}_{\rm HIGH}}$ then
18:           ON $\leftarrow\mathbf{\rm false}$triangleright switch-off
19:        else if ${\cal L}_{\lambda/E}<\gamma_{{}_{\rm LOW}}$ then
20:           ON $\leftarrow\mathbf{\rm true}$triangleright switch-on
21:        end if
22:    end while
23:    $\widehat{\mathbf{H}}_{b}\leftarrow{\rm predict}(\mathbf{H}_{b},\mathbf{P},{\rm net},\phi)$triangleright inference on band $b$
24:end for
25:return $\widehat{\mathbf{H}}$
Algorithm 1 $\rho$-PNN pseudo-code.
@@ IV-A2 Limiting the Number of Iterations
$N^{\rm max}_{b}=N_{0}+\Delta N(1-\mathbf{c}_{b})$ (8)
with $c_{b}$ the correlation coefficient between bands $b$ and $b-1$. Therefore, the maximum number of iterations is $N_{0}\geq N_{S}^{\rm max}$ in case of perfect correlation ($\mathbf{c}_{b}=1$) and $N_{0}+\Delta N$ in case of orthogonal bands ($\mathbf{c}_{b}=0$). With this rule, we can easily set an overall bound to the total number of iterations $N$ as follows
@@
$\Delta N=\frac{\eta BN_{0}}{B-\sum_{b}\mathbf{c}_{b}}$ (10)
Fig. 9 shows an example of how iterations distribute over the various bands of a PRISMA image. Here we set $N_{S}^{\rm max}=20,N_{0}=50$ and $\eta=8$. $N_{b|S}$ is shown in yellow and $N_{b}$ in brown, with its upper bound $N^{\rm max}_{b}$ in light gray. Although a maximum of $(1+\eta)N_{0}=450$ iterations per band is allowed on average, only 47 are used on the average. Most of the times, tuning stops because $N_{b|S}$ reaches its maximum, $N_{S}^{\rm max}=20$. The total number of iterations rarely exceeds $N_{0}=50$ and only for a group of bands, on the boundary between the visible and NIR ranges, the $N_{b}=N_{b}^{\rm max}$ exit condition is activated.
@@
$\gamma_{{}_{\rm HIGH}}$ 0.65 Upper limit for ${\cal L}_{\lambda/E}$
$\gamma_{{}_{\rm LOW}}$ 0.59 Lower limit for ${\cal L}_{\lambda/E}$
$\beta_{0}$ 2 Initial $\beta$, weight of the spatial loss term
$\epsilon$ 0.007 Growth factor of ${\cal L}_{\lambda}$ for $\beta$ adaptation
$N_{S}^{\rm max}$ 20 Target # of tuning iter.s with active spatial loss
$N_{0}$ 80 Max # of tuning iter.s when $\mathbf{c}_{b}=1$
$\eta$ 30 Global param. controlling total # of tuning iter.s
@@ IV-B Assessment Methodology
To assess the quality of the proposed solution we take advantage of the Benchmarking toolbox [40] which provides a publicly available set of rich and diverse PAN-HS PRISMA images, a large set of SotA comparative methods, and several performance evaluation tools based on the most credited assessment indexes.
TABLE IV: PAN-MS PRISMA benchmark datasets for testing.
Date Location Land covers
05/09/22 Cagliari, Italy Roofs, Streets, Crops, Hills, Water
24/08/23 Udine, Italy Roofs, Streets, Crops, Water
08/09/23 Ford County, Kansas Streets, Crops
20/11/23 Macuspana, Tabasco Streets, Crops
Real FR test image size: 1200$\times$1200
Resolution ratio $R=6$
Bands $B=159/239$
Resolution: 30 m (RR) or 5 m (FR)
PRISMA satellites are equipped with a sensor operating in the 400–700 nm spectral range that provides a panchromatic image with 5 m spatial resolution, and two hyperspectral sensors that acquire 66 bands at VNIR wavelengths, 400–1010 nm, and 173 bands at SWIR wavelengths, 920–2505 nm, with 30 m resolution. The dataset released by [40] includes 16 images (12 for training and validation, 4 for testing) comprising 159 spectral bands after removing noisy and uninformative channels. For our experiments we use the four test images of the Toolbox, see Tab. IV. From each original PRISMA PAN-HS pair (PAN/HS size: 3600$\times$3600/600$\times$600) two test datasets were created: a FR one (PAN/HS size: 1200$\times$1200/200$\times$200) obtained by cropping; a RR one (PAN/HS/GT size 600$\times$600/100$\times$100/600$\times$600) obtained by resolution downgrading of the whole PAN-HS pair, with the original FR HS playing the role of GT.
Tab. V gathers the comparative methods which can be roughly grouped in four categories: component substitution (CS), multiresolution analysis (MRA), model-based optimization (MBO), and deep learning (DL). The interested Reader is referred to [40] for details on these methods.
TABLE V: HS pansharpening methods.
@@
The assessment follows the well-established Wald’s protocol [82] which considers two contexts: a reduced resolution framework, where suitably degraded (in resolution) datasets are used for assessment with GT (where the ground truth is given by the original full resolution HS image), and a full resolution framework for assessment without GT but using the original data as input. Accordingly, the quality indexes, gathered in Tab. VI, are split in two groups for reduced resolution (RR) and full resolution (FR) assessment.
@@
TABLE VI: HS pansharpening quality assessment indexes.